Circular No. 8623

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN C. W. Hergenrother, Lunar and Planetary Laboratory, reports his re-

covery of this comet on images taken with the 1.2-m reflector (+ Gunn r filter) at Mount Hopkins in bright moonlight, the observations below apparently belonging to component C. The comet is strongly condensed with a coma 6′′ in diameter and a short fan tail 8′′ long in p.a. 300^{o}.

2005 UT  α_{2000}  δ_{2000}  Mag.
Oct. 22.49350  10^{h}52^{m}53.^{s}72  +16^{o}59′35′′.6
22.50150  10 52 54.44  +16 59 31.9
22.50970  10 52 55.14  +16 59 28.7
22.51780  10 52 55.82  +16 59 25.8  19.3
24.49490  10 55 48.12  +16 46 51.6
24.50334  10 55 48.76  +16 46 46.9
24.51130  10 55 49.46  +16 46 45.8  19.2

The following orbital elements by B. G. Marsden are from 224 observations, 1995–2005 (mean residual 0′′.7; nongravitational parameters A_{1} = +1.33, A_{2} = −0.0520). The correction to the prediction by S. Nakano (MPC 48382; 2005 Comet Handbook) is ΔT = −0.43 day. Component C will pass 0.08 AU from the earth on 2006 May 12. Predictions for the components seen at previous returns have been made recently by Sekanina (CBET 237; ICQ 27, 225).

Epoch = 2006 May 25.0 TT

T = 2006 June 6.9497 TT  ω = 198.^{o}8039
e = 0.693192  Ω = 69.8955 }2000.0
q = 0.939135 AU  i = 11.3960

a = 3.060989 AU n^{o} = 0.1840396 P = 5.355 years

SUPERNOVA 2005hg IN UGC 1394 Further to IAUC 8616, K. Shimasaki and W. Li report the LOSS dis-

covery of an apparent supernova on unfiltered KAIT images taken on Oct. 25.27 (at mag 18.6) and 26.25 UT (mag 18.5). SN 2005hg is located at α = 1^{h}55^{m}41^{s}.87, δ = +46^{o}47′47′′.4 (equinox 2000.0), which is 3′′.7 west and 19′′.6 south of the nucleus of UGC 1394. KAIT images taken on Oct. 12.25 (limiting mag 20.0) and 20.24 (limiting mag 19.5) showed nothing at this position. SN 2002gc also appeared in UGC 1394.

2005 October 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green